OpenStudy (anonymous):

A right isosceles triangular plate is vertically submerged below the surface of a fluid of weight density w. The top of the plate is at the surface of the fluid. Find the fluid force on the plate in terms of w.

OpenStudy (anonymous):



OpenStudy (anonymous):

Uhm.... .-.

OpenStudy (anonymous):

I think the fluid force is then: \[F = w \int\limits_{a}^{b} yf(y)~~ dy\] is this right? :3

OpenStudy (anonymous):

Yes? :3 haha /.\

OpenStudy (anonymous):

yeah I'd say this is physics and not math

OpenStudy (anonymous):

It's for my math class though xD

OpenStudy (anonymous):

ah calc

OpenStudy (anonymous):

I think carrying on from what I payed attention to in class the limits for a and b would be like: a = 0 , b= 3 and f(y) = y so: \[F = w \int\limits_{0}^{3} y(y) ~~dy\] \[w \int\limits_{0}^{3} y^2~ dy\] \[w * \left[ \frac{ 1 }{ 3 }y^3 \right]^3_0\] \[F = 9w\] ?

OpenStudy (wolfe8):

If I'm not going way out there, isn't that formula basically the density times the area? Hmm how did you get f(y)=y?

OpenStudy (anonymous):

f(y) = y because it is the length of the horizontal part from the water surface, if that isn't wrong..

OpenStudy (anonymous):

the length of the horizontal part is at a depth of y from the water surface* sorry getting sleepy >.<

OpenStudy (wolfe8):

http://tutorial.math.lamar.edu/Classes/CalcII/HydrostaticPressure.aspx I think you need to multiply by gravity as well.

OpenStudy (anonymous):

The formula involving gravity is for hydrostatic pressure not force xD F = (w*h)*A is the formula for fluid force..

OpenStudy (anonymous):

I'll just call up my friend and ask her what she got xD thanks for helping anyways :)

OpenStudy (wolfe8):

Dang that's what I've had in mind all this time. I think it looks right to me.